You earn $74.25 for working 9 hours. How much will you earn for working 15 hours?
Answer by stanbon(75887)

You earn $74.25 for working 9 hours.
How much will you earn for working 15 hours?
--------------------------
Set up a proportion and solve for "x":
x/15 = 74.25/9
----
x = 15(74.25/9)
----
x = $123.75
=================
Cheers,
Stan H.